OpenStudy (anonymous):

show me how to do 6n+18-12n= -36

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

18-6n=-36

OpenStudy (anonymous):

6n -12n + 18 = -36 -6n = -36 -18 -6n = -54 n = 9

OpenStudy (anonymous):

thanks! :D

OpenStudy (anonymous):

what about this one

OpenStudy (anonymous):

n/7 - 14 = 21

OpenStudy (anonymous):

i would multiply by 7 to remove the fraction (i just do not like fractions)

OpenStudy (anonymous):

oh nvm i figured that one out

OpenStudy (anonymous):

actually i meant his one 41= - n/8 + 54

OpenStudy (anonymous):